What is the interference in Wm/2 of 85dB sound?


1
Expert's answer
2021-07-28T03:34:03-0400

The intensity level of the sound

"\beta=10\lg\frac{I}{I_0}"

So, the intensity of the sound

"I=I_0*10^{\beta/10}\\=10^{-12}*10^{85/10}=3.2*10^{-4}\:\rm W/m^2"